OpenStudy (anonymous):

A game has 40 chips, each labeled with a number from 1 through 40. Lisa randomly chooses one chip. What is the probability that she chooses 25 or a multiple of 10? A.1/400 B.1/80 C.1/10 D. 1/8

OpenStudy (anonymous):

1/40 x 4/40 = 1/8

OpenStudy (anonymous):

Wrong equation, right answer. There are 5 options which suit the criteria: 10, 20, 25, 30, and 40. 5/40=1/8.

OpenStudy (anonymous):

darn, I never get it rightp

OpenStudy (anonymous):

You got the right answer. You know what, I think you just had X instead of +.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol okay then xD